40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog, Page 3 of 9  Not logged in ELOG logo
ID Date Author Type Category Subject
  311   Thu Jan 10 20:42:54 2019 KojiOpticsCharacterizationFSR / HOM Test of OMC SN002

OMC SN002 = Former LHO OMC which CM1 was destroyed by the lock loss pulse in 2016. This OMC needs to be optically tested before storage.

The test items:

  • [done] FSR measurement with offset PDH locking (FM->AM conversion)
  • [done] FSR/finesse measurement with the EOM RFAM injection
  • [done] TMS measurement with input miaslignment and the trans RFPD misalignment: with no PZT offset
  • [done] TMS measurement with input miaslignment and the trans RFPD misalignment: with PZT offsets
     
  • PZT response
  • Mirror cleaning
  • Power budget
  • Diode alignment: shim height
  • PD/QPD alignment
  310   Thu Nov 1 19:57:32 2018 AaronOpticsGeneralMontecarlo simulation of the phase difference between P and S pols for a modeled HR mirror

I'm still not satisfied/done with the solution to this, but this has gone too long without an update and anyway probably someone else will have a direction to take it that prevents me spinning my wheels on solved or basic questions.

The story will have to wait to be on the elog, but I've put it in the jupyter notebook. Basically:

  • I considered the polarization-separated OMC in several configurations. I have plots of DARM referred noise (measured free-running and controlled noise for the current OMC, thermal theoretical noise curve, scattered light) for the case of such an OMC with one lambda/2 waveplate oriented at 45 degrees. This is the base case.
  • I also considered such an OMC with a lambda/2 both before and after the OMC, where their respective polarization axes can be arbitrary (I look at parameter space near the previous case's values).
    • I optimize the BHD angle to balance the homodyne (minimize the E_LO^2 term in the homodyne readout).
    • I then optimize the rotations of the lambda/2 polarization axes to minimize the noise
    • For the optimum that is closest to the base case, I also plotted DARM referred length noise.

 

It's clear to me that there is a way to optimize the OMC, but the normalization of my DARM referred noise is clearly wrong, because I'm finding that the input-referred noise is at least 4e-11 m/rt(Hz). This seems too large to believe. 

Indeed, I was finding the noise in the wrong way, in a pretty basic mistake. I’m glad I found it I guess. I’ll post some plots and update the git tomorrow. 

  309   Thu Sep 27 20:19:15 2018 AaronOpticsGeneralMontecarlo simulation of the phase difference between P and S pols for a modeled HR mirror

I started some analytic calculations of how OMC mirror motion would add to the noise in the BHD. I want to make some prettier plots, and am adding the interferometer so I can also compute the noise due to backscatter into the IFO. However, since I've pushed the notebook I wanted to post an update. Here's the location in the repo.

I used Koji's soft limit of 0.02 degrees additional phase accumulation per reflection for p polarization.

  308   Sun Sep 23 19:42:21 2018 KojiOpticsGeneralMontecarlo simulation of the phase difference between P and S pols for a modeled HR mirror

[Koji Gautam]


With Gautam's help, I ran a coating design code for an HR mirror with the standard quarter-wave design. The design used here has 17 pairs of lambda/4 layers of SiO2 and Ta2O5 (=34 layers) with the fused silica as the substrate to realize the transmission of tens of ppm. At the AOI (angle of incidence) of 4 deg (=nominal angle for the aLIGO OMC), there is no significant change in the reflectivity (transmissivity). With 95% of the case, the phase difference at the AOI of 4 deg is smaller than 0.02 deg for given 1% fluctuation (normal distribution) of the layer design and the refractive indeces of the materials. Considering the number of the OMC mirrors (i.e. 4), the total phase shift between P and S pols is less than 0.08 deg. This makes P and S resonances matched well within 1/10 of the cavity resonant width (360/F=0.9deg, F: Finesse=400).

Of course, we don't know how much layer-thickness fluctuation we actually have. Therefore, we should check the actual cavity resonance center of the OMC cavity for the polarizations.

Attachment 1 shows the complex reflectivity of the mirror for P and S pols between AOIs of 0 deg and 45 deg. Below 30 deg there is no significant difference. (We need to look at the transmission and the phase difference)

Attachment 2 shows the power transmissivity of the mirror for P and S pols between AOIs of 0 deg and 45 deg. For the purpose to check the robustness of the reflectivity, random fluctuations (normal distribution, sigma = 1%) were applied to the thicknesses of each layer, and the refractive indices of Silica and Tantala. The blue and red bands show the regions that the 90% of the samples fell in for P and S pols, respectively. There are median curves on the plot, but they are not well visible as they match with the ideal case. This figure indicates that the model coating well represents the mirror with the transmissivity better than 70ppm.

Attachment 3 shows the phase difference of the mirror complex reflectivity for P and S pols between AOIs of 0deg and 45deg. In the ideal case, the phase difference at the AOI of 4deg is 1x10-5 deg. The Monte-Carlo test shows that the range of the phase for 90% of the case fell into the range between 5x10-4 deg and 0.02 deg. The median was turned to be 5x10-3 deg.

Attachment 4 shows the histogram of the phase difference at the AOI of 4deg. The phase difference tends to concentrate at the side of the smaller angle.

Attachment 1: reflectivities.png
reflectivities.png
Attachment 2: transmission.png
transmission.png
Attachment 3: phase_diff.png
phase_diff.png
Attachment 4: phase_difference_histogram.png
phase_difference_histogram.png
  307   Wed Aug 29 11:06:30 2018 KojiGeneralGeneralRF AM RIN and dBc conversion

0. If you have an RF signal whose waveform is 1 \times \sin(2 \pi f t), the amplitude is constant and 1.

1. If the waveform [1+0.1 \sin(2 \pi f_{\rm m} t)] \sin(2 \pi f t), the amplitude has the DC value of 1 and AM with the amplitude of 0.1 (i.e. swing is from 0.9 to 1.1). Therefore the RMS RIN of this signal is 0.1/1/Sqrt(2).

2. The above waveform can be expanded by the exponentials.

\left[-\frac{1}{2} i e^{i\,2\,\pi f t} + 0.025 e^{i\,2\,\pi (f-f_{\rm m}) t}- 0.025 e^{i\,2\,\pi (f+f_{\rm m}) t} \right] - {\rm C.C.}

Therefore the sideband carrier ratio R is 0.025/0.5 = 0.05. This corresponds to 20 log10(0.05) = -26dBc


In total, we get the relationship of dBc and RIN as {\rm dBc} = 20 \log_{10}(\rm{RIN}/\sqrt{2}), or R = RIN/sqrt(2)

  306   Thu Aug 9 11:24:29 2018 KojiGeneralCharacterizationModulation Index Test Setup at 40m Lab

[Rich Koji]

The impedances of the new LLO EOM were measured with the beat note setup at the 40m PSL (as described in the previous ELOG entry.

At the target frequencies (9.1MHz, 24.1MHz, 45.5MHz, 118.3MHz), the modulation responses were (0.09, 2.9e-3, 0.053, 0.021) rad/V.

This corresponds to the requirement for the driving power as follows.

Frequency
[MHz]
Response
[rad/V]
modulation depth 
required (LHO) [rad]
Required
drive [Vpk]
Required
drive [dBm]
    9.1 0.09 0.22 2.4 17.8
  24.1 2.9e-3   0.014 4.8 23.7
  45.5   0.053 0.28 5.3 24.5
118.3   0.021   0.010   0.48   3.6

 

Attachment 1: modulation_depth.pdf
modulation_depth.pdf
Attachment 2: modulation_depth_zoom.pdf
modulation_depth_zoom.pdf
  305   Wed Aug 8 17:32:56 2018 Rich AbbottGeneralCharacterizationModulation Index Test Setup at 40m Lab

Attached is a block diagram of the test setup used in the 40m lab to measure the modulation index of the IO modulator

Attachment 1: 40mLabModIndexSetup.pdf
40mLabModIndexSetup.pdf
  304   Tue Aug 7 15:43:12 2018 KojiElectronicsCharacterizationNew LLO EOM stuffed

[Rich, Dean, Koji]

Stuffed all inductors for the new LLO EOM. As the impedances were sensitive to the positions of the inductors in the housing, they were glued with a glue gun.
Also the lid of the housing significantly change the stray capacitance and lowers the resonant frequency (meaning lowers the Q too), we decided to tune the matching circuit without the lid.

The attached plots show the measured impedances. They all look well tuned and matched. We will prepare and perform the optical measurement at the 40m.

Attachment 1: P_20180806_154457.jpg
P_20180806_154457.jpg
Attachment 2: impedance_eom.pdf
impedance_eom.pdf
Attachment 3: impedance_eom_zoom.pdf
impedance_eom_zoom.pdf
  303   Thu Jul 26 20:57:07 2018 KojiElectronicsCharacterization9MHz port tuned impedance

[Rich Koji]

The 9MHz port was tuned and the impedance was measured.

Attachment 1: impedance_eom.pdf
impedance_eom.pdf
Attachment 2: eom9.pdf
eom9.pdf
Attachment 3: eom_models_9MHz.pdf
eom_models_9MHz.pdf
  302   Wed Jul 4 18:30:51 2018 KojiElectronicsCharacterizationEOM circuit models

The circuit models for the 3IFO EOM (before mods) were made using LISO.
Then the modification plan was made to make it a new LLO EOM.

Impedance data, LISO model, Mathematica files are zipped and attached at the end.

Attachment 1: eom_models.pdf
eom_models.pdf eom_models.pdf eom_models.pdf
Attachment 2: eom9.pdf
eom9.pdf
Attachment 3: eom24.pdf
eom24.pdf
Attachment 4: eom45.pdf
eom45.pdf
Attachment 5: 180704_3IFO_EOM_model.zip
  301   Tue Jul 3 12:07:47 2018 Rich AbbottElectronicsCharacterizationNotes on 3rd IFO EOM

Attached please see my notes summarizing the models for the electrodes and inductors within the 3rd IFO EOM

Attachment 1: EOM_Analysis2.pdf
EOM_Analysis2.pdf EOM_Analysis2.pdf
  300   Mon Jul 2 15:27:31 2018 Rich AbbottElectronicsGeneralWork on EOM (3rd IFO unit)

Koji, Rich

We took apart the unit removed from the 3rd IFO (Unit serial number aLIGO #3, XTAL 10252004) to see what makes it tick.  Koji has done a fine job of adding the plots of the impedance data to this log book.  Attached are some details of the physical construction showing the capacitor values used in shunt before the coils.

Attachment 1: EOM3_aLIGO_3rdIfo.JPG
EOM3_aLIGO_3rdIfo.JPG
  299   Mon Jul 2 12:29:01 2018 KojiElectronicsCharacterizationImpedances of individual components (3IFO EOM)

[Rich Koji]

The impedances of the individual components from the 3IFO EOM (before modification) were tested.
Each component was modeled by LISO. The LISO model (in PDF and txt) are attached at the end of the entry.

Coils
There are three inductors taken from the EOM unit. They showed the Q ranging from 150~300.
Their impedances are compared with the coil taken from the 9MHz port of the spare EOM (=current LHO EOM).
The inductance of the 8.7MHz inductor indicated higher L but still higher Q.

Todd made a replica of the 45.3MHz coil. He used a silver plated wire and it actually showed highest Q of ~400.

Crystal capacitance
The crystal capacitances were measured by attaching a test rig on the DB15 connector of the crystal housing. The rig was calibrated such that the impedances of the attched components on the rig were measured. They showed somewhat similar feature with parasitic resonances at ~50MHz. Above this frequnecy the capacitance went down (i.e. Abs(Z) went up). This indicates there are stray series LCR in pararrel to the crystal. Not sure what is the cause of this.

The central (24.1MHz) port showed smaller capacitance. This probably means the plates for the central port is smaller. Not sure the actual dimensions of the plates for this unit.
 

Attachment 1: impedance_coils.pdf
impedance_coils.pdf
Attachment 2: impedance_xtals.pdf
impedance_xtals.pdf
Attachment 3: q_coils.pdf
q_coils.pdf
Attachment 4: component_models.pdf
component_models.pdf component_models.pdf component_models.pdf component_models.pdf component_models.pdf component_models.pdf component_models.pdf component_models.pdf
Attachment 5: liso_models.zip
  298   Mon Jul 2 11:30:22 2018 KojiElectronicsCharacterization3IFO EOM impedance measurement

[Rich Koji]

3IFO EOM (before any modification) was tested to measure the impedance of each port.

The impedance plot and the impedance data (triplets of freq, reZ, imZ) were attached to this entry.

Attachment 1: impedance_eom.pdf
impedance_eom.pdf
Attachment 2: EOM_Z_DATA.zip
  297   Wed May 30 17:44:23 2018 KojiOpticsCharacterization3IFO EOM surface check

3IFO EOM dark microscope images courtesy by GariLynn and Rich

Attachment1/2: Hole #1
Attachment3/4: Hole #2
Attachment5: Hole #2

Attachment 1: IMG_5756.JPG
IMG_5756.JPG
Attachment 2: IMG_5757.JPG
IMG_5757.JPG
Attachment 3: IMG_5758.JPG
IMG_5758.JPG
Attachment 4: IMG_5759.JPG
IMG_5759.JPG
Attachment 5: Looking_at_Hole2.png
Looking_at_Hole2.png
  296   Wed May 30 16:40:38 2018 KojiMechanicsCharacterizationEOM mount stability test

https://awiki.ligo-wa.caltech.edu/wiki/EOM_Mount_Stability

  295   Tue May 15 19:53:45 2018 KojiOpticsGeneralEOM Q comparison

Qs' were estimated with a lorentzian function (eye fit)

aaa

Current LHO EOM (final version, modulation depth measurement 2018/4/5)
f0=9.1MHz, Q=18
f0=45.38MHz, Q=46
f0=118.05MHz, Q=30

Prev LHO EOM (RF transmission measurement 2018/4/13)
f0=9.14MHz, Q=53
f0=24.25MHz, Q=55
f0=45.565MHz, Q=62;

3IFO EOM (RF transmission measurement 2018/4/23)
f0=8.627MHz, Q=53
f0=24.075MHz, Q=60
f0=43.5MHz, Q=65

  294   Sat May 5 22:51:04 2018 KojiOpticsGeneral3IFO EOM Optical test

The 3IFO EOM test performed at the 40m. Result: 40m ELOG 13819

  293   Thu May 3 21:45:58 2018 awadeGeneralLoan / LendingBorrowed toaster oven

I’ve borrowed the black and decker toaster oven to dry some sonicated parts. It is temporarly located in the QIL lab. 

Attachment 1: 9CE80545-7A58-4236-B7E3-1EE6C4042DAA.jpeg
9CE80545-7A58-4236-B7E3-1EE6C4042DAA.jpeg
  292   Mon Apr 2 17:27:04 2018 KojiOpticsCharacterizationaLIGO EOM test

2nd optical test http://nodus.ligo.caltech.edu:8080/40m/13725

  291   Thu Feb 22 20:21:02 2018 KojiOpticsCharacterizationaLIGO EOM test

POSTED to 40m ELOG

  290   Thu Nov 30 12:18:41 2017 StephenGeneralGeneralPreparation for Modal Testing on 4 December

Norna Robertson, Stephen Appert ||  29 Nov 2017, 2 pm to 4 pm  ||  227 Downs, CIT

We made some preparations for modal testing, but did not have enough time to make measurements. Below is an after-the-fact log, including some observations and photos of the current state of the OMC bench.

  1. Previous testing results at T1700471 (technical note in progress as of 30 Nov 2017).
    1. One goal of the next round: add damping material to equate with damping material of T1600494.
    2. Second goal of the next round: use a more localized sweep to better resolve the body mode around 1080 Hz -1100 Hz
  2. Transport Fixture was opened without issue, revealing the "Top" (suspending and cable routing) surface of the bench. Damping stacks were still in place from previous testing
  3. We removed the bolts from the damper stacks, but found that all masses with metal-viton interfaces had adhered to viton washers, causing the stacks to stick together.
    1. By using an allen key as a lever to wedge apart bottom mass and the bracket where they were joined by a viton washer, we separated the masses from the bracket.
    2. An allen key was used as a lever to push apart the two masses, which were also joined by a viton washer
    3. Once exposed, viton washers were pried from metal surfaces.
  4. After the damper stacks had been detached from the  No viton washer appeared to leave any residue or particulate - the separated parts all appeared as clean as they had been at the onset.
  289   Mon Nov 27 20:24:24 2017 KojiGeneralGeneralA former LHO PD (Trans) removed from the OMC #002 for the shipment to Stockholm

Attachment 1: The PD was removed from the transmission side of the OMC #002 (former LHO OMC - the one blasted by the optical pulse in Aug 2016).
It was confirmed that the PD has the scribing mark saying "A".

Attachment 2: This diode had no glass cap on it. The photodiode sensitive element is still intact. For ease of handling, it should be kept in a cage. There are four cages in the OMC lab, but they are ocuppied with the High QE PDs and others. So, the cage for this PD was offered by Rich from his office, meaning the cage was not clean.

Attachment 3: The sensor side is capped by a plate. This cap can be removed by unscrewing the two cap screws in the photo.

Attachment 4: The PD legs are shorted. (Just to match the style with the LLO one).

Attachment 5: Wrapped with AL foil and double bagged. (Repeat: It is not anything clean.)

Attachment 6: The bag was left on Rich's desk.

Attachment 1: IMG_2826.JPG
IMG_2826.JPG
Attachment 2: IMG_2835.JPG
IMG_2835.JPG
Attachment 3: IMG_2833.JPG
IMG_2833.JPG
Attachment 4: IMG_2836.JPG
IMG_2836.JPG
Attachment 5: IMG_2837.JPG
IMG_2837.JPG
Attachment 6: DSC_0546.JPG
DSC_0546.JPG
  288   Fri Sep 8 15:14:05 2017 KojiFacilityGeneralPreparation for the plumbing work

[Steve, Aaron, Koji]

We've finished the preparation for the forthcoming plumbing work on (nominally) Sept 16th Saturday.
We've covered most of the west side of the OMC lab with plastic sheets and wraps.

Some tips:

  • The plastic sheets Eric gave us were a bit too thin and pron to got torn. Thicker sheets are preferable.
  • The blue tape that Eric gave us was very useful.
  • The stretch wrap film, which I bought long time ago, was so useful. Office Depot "Office Depot(R) Brand Stretch Wrap Film, 20 x 1000 Roll, Clear" PN: 445013
  • We also used patches of Kitchen Trash Bags to cover some small opening of the large sheets. Office Depot "Glad(R) Tall Kitchen Trash Bags, 13 Gallon, White, Box Of 28" PN 269268
Attachment 1: DSC_0405.jpg
DSC_0405.jpg
Attachment 2: DSC_0406.jpg
DSC_0406.jpg
Attachment 3: DSC_0407.jpg
DSC_0407.jpg
Attachment 4: DSC_0408.jpg
DSC_0408.jpg
Attachment 5: DSC_0409.jpg
DSC_0409.jpg
Attachment 6: DSC_0410.jpg
DSC_0410.jpg
Attachment 7: DSC_0411.jpg
DSC_0411.jpg
  287   Sat Jul 29 21:42:51 2017 KojiElectronicsCharacterizationPDH amp

The polarities indicated in the right circuits were opposite, obviously.

  286   Sat Jul 29 18:44:38 2017 ranaElectronicsCharacterizationPDH amp

attachment 6: DCPD preamp looks like the opamp is wired for positive feedback?

  285   Wed Jul 5 16:59:44 2017 KojiGeneralGeneralThe OMC #002 was packed

[Stephen Koji]

The OMC #002 was packed for the transportation to Downs.

===> And transported to Downs 227 on Jul 6th.

Attachment 1: DSC_0360.JPG
DSC_0360.JPG
  284   Sat Jul 1 21:33:18 2017 KojiGeneralGeneralSome purchase notes

- Forgot to close the cylinder valve...

v HEPA prefilter (20"x20"x1" MERV 7)

- Replace the filter for the air conditioning

v Texwipe TX715 SWAB http://www.texwipe.com/store/p-817-tx715.aspx

v Gloves ~3 bags
VWR GLOVE ACCTCH NR-LTX SZ7.0 PK25 79999-304 x3
VWR GLOVE ACCTCH NR-LTX SZ7.5 PK25 79999-306 x1

v Vectra IPA soaked cloths

v Sticky mats


  • GLOVE ACCTCH NR-LTX SZ7 PK25 / 79999-304 / PK4
  • GLOVE ACCTCH NR-LTX SZ7.5 PK25 / 79999-306 / PK1
  • WIPER 100% IPA 23X23CM PK50. / TWTX8410 / PK2
  • SWAB CLEANTIPS ALPHA PK100. / TW-TX715 / PK1
  • MAT CLEAN ROOM 18X36IN BLUE / 89021-748 / CS1 (Qty4)
  • FILTER PLET AIR MERV8 20X20X1 / 78002-422 / EA4 / Direct from Supplier

ORDERED AUG 9, 2017

  283   Sat Jul 1 15:29:57 2017 KojiOpticsGeneralBlack glass cleaning / Final bonding for the emergency repair for OMC #002

[Alena, Koji]

Report of the work on June 30.

1. Cleaning of the black glass beam dumps

As reported in the previous entry, the beam dumps on the OMC breadboard exhibited accumulation of dusts or contaminants on the black glass surfaces. We worried about transfer of the dusts over a period or of the contaminant during baking. It was already known that the contaminants are persistent and not easy to remove only by drag wiping with IPA. So Alena brought a set fo tools to try. Here is the procedure described.

- Inventory (Attachment 1): A small glass beaker, TX715 Alpha® Sampling Swab, plastic brushes, syringes with pure IPA, inspection flash light, Vectra IPA soaked wipes

- Apply clean IPA on a brush. Some IPA should be removed by the IPA soaked wipe so as not to splash IPA everywhere. Rub a glass surface with the brush while the surface is inspected by the flash light. The strokes migrate the contaminants to the direction of wiping. So the brush should be moved outward. This does some cleaning, but it is not enough to remove smudges on the surface. Occasionally clean the brush with IPA poured in the small beaker.

- Apply clean IPA on a swab. Rub the surface with the swab outward. This removes most of the visible smudges.

We decided not to apply FirstContact on the beam dumps at this occasion. In any case, we need to apply FC on all the optical surfaces after the baking. We judged that the current cleanliness level of the beam dump does not affect the over all contamination of the OMC considering the FC application after the baking.

2. Gluing of the reinforcement Al bars on the delaminated Invar mounting brackets
One of the mounting bracket (=invar shim) on the top side (= suspension I/F side) showed the sign of delamination (Attachment 3). This invar is the one at the beam entrance side (Attachment 2).

EP30-2 was mixed as usual: 6g of EP30-2 was mixed with 0.3g glass sphere. The glue was tested with a cooking oven and the result was perfect. The glue was applied to two Al bars and the bars were attached on the long sides of the invar shim with the beveled corner down (to avoid stepping on the existing original epoxy) (Attachments 4, 5). The photo quality by my phone was not great. I will take better photos with a better camera next week.


Glue condition was checked on Monday Jul 3rd. It was all good. New photos were taken. OMC #002 Repair - Gluing of reinforcement AL bars

Attachment 1: DSC_0347.jpg
DSC_0347.jpg
Attachment 2: DSC_0348.jpg
DSC_0348.jpg
Attachment 3: DSC_0350.jpg
DSC_0350.jpg
Attachment 4: DSC_0352.jpg
DSC_0352.jpg
Attachment 5: DSC_0354.jpg
DSC_0354.jpg
  282   Fri Jun 23 10:55:07 2017 KojiOpticsGeneralDust layer on black glass beam dumps?

I wondered why the black glass beam dumps looked not as shiny as before. It was in fact a layer of dusts (or contaminants) accumulated on the surface.
The top part of the internal surface of the black glass was touched by a piece of lens tissue with IPA. The outer surface was already cleaned. IPA did not work well i.e. Required multiple times of wiping. I tried FirstContact on one of the outer surface and it efficiently worked. So I think the internal surfaces need to be cleaned with FC.

Attachment 1: black_glass_dust.JPG
black_glass_dust.JPG
  281   Fri Jun 23 01:58:11 2017 KojiOpticsGeneralOMC #002 Repair - CM1 gluing

[Alena, Koji]

Jun 21: Alena and Koji worked on gluing of the CM1 mirror on the OMC breadboard #002. This is an irregular procedure. Usually, the PZT mirror subassembly is prepared before the mounting prism is glued on the breadboard. In this occasion, however, a PZT and a mirror are bonded on an existing prism because only the damaged mirror and still functional PZT were debonded from the mouting prism.

For this purpose, the mirror and the PZT were fixed on the mounting prism with the modified fixture set (D1600338). The original PZT was reused, and the new mirror #8 was used. The alignment of the mirror was checked OK using the cavity beam before any glue was applied. The arrow of the CM mirror is facing up.

We mixed 8g EP30-2 (it was almost like 3~4 pushes) and 0.4g glass sphere bond lining. Along with EP30-2 procedure, the bond was mixed in an Al pot and tested with 200degF (~93degC) preheated the oven for 15min. The cured bond showed perfect dryness and crispness. The bond was painted on the PZT and the PZT was placed on the fixture. Then more bond was painted on the other side of the PZT. The mirror was placed in the fixture. The spring-loaded front plate was fixed, and the breadboard was left for a day. (Attachment 1~3)

Jun 22: The fixture was removed without causing any visible delamination or void. The attachment 4~6 show how wet the joint is (before baking). There were some excess of EP30-2, which bonded the fixture and the mounting prism as usual. The fixture was detached by prying the front piece against the rear piece with a thin allen key. Some of the excess bond on the mounting prism was removed by scratching.

The alignment of the cavity was checked with the cavity beam and it is still fine.

More photos can be found here: Link to Google Photos Album "OMC #002 Repair - CM1 gluing"

Attachment 1: IMG_0857.JPG
IMG_0857.JPG
Attachment 2: IMG_0859.JPG
IMG_0859.JPG
Attachment 3: IMG_0860.JPG
IMG_0860.JPG
Attachment 4: IMG_0865.JPG
IMG_0865.JPG
Attachment 5: IMG_0868.JPG
IMG_0868.JPG
Attachment 6: IMG_0864.JPG
IMG_0864.JPG
  280   Tue Jun 6 22:00:36 2017 KojiGeneralConfigurationTrans RF PD setup

- Replaced the PZT with the one used from the beginning. This must be PZT #21. After the replacement, the spot positions look very good. I even went up. So I decided this is the configuration to proceed to the gluing. The CM1 mirror has the HR arrow at the top.

- The input beam was realigned w.r.t. the OMC.

- Tried to use the IR viewer with the new rechargable battery brought from the 40m. But the view still didn't work. The possibility is a) the viewer is broken b) the battery is empty.

- Tried to use the stainless clean regulartor for the UHP N2. The outlet has a short tube with a different diameter. The O.D. of the old tube is 6.3mm, while the new one is 9.5mm. If I insert the thinner tube in the new tube, it approximately fits. But I don't believe this is the way...

  279   Tue Jun 6 00:49:48 2017 KojiGeneralConfigurationTrans RF PD setup

Last week, I further worked on the RF system to install 20dB coupler on the agilent unit and setup the R channel. This allowed me to make the FSR/TMS measurement of the OMC.

And today several optical improvement has been done.

- The input/output fiber couplers were adjusted to have the maximum transmission through the PBS right before the OMC.
- The HWP on the output side of the faraday was adjusted to have ~40mW input to the OMC.

Then, the OMC curved mirror is now held by the new in-situ gluing fixture instead of the conventional fixture attached upside down.
The OMC was ocked again and the input alignment was adjutsed. The fixture is blocking the QPD path, so it's not possible to confirm the proper alignment of the cavity (w.r.t. the QPD paths).

The precise positions of the spots could not be confirmed as the battery of the IR viewer was empty. Quick check of the spots by the card tells that the spot on the CM2 (PD side) is slightly too close to FM2 (output coupler). I wonder if this could be solved by rotating the curved mirror.

Otherwise everything look good. Let's try to glue the curved mirror tomorrow.

Note: Spot on CM2 is too close to the edge of the hole on the mounting prism. The meausrementof CM1 is telling that the curverture center is located 2.7mm upper side of the center of the mirror if the HR side arrow is up (and it is the case). If we move the arrow to the QPD path side (90deg CW viewed from the face side), this corresponds to ~1.1mrad CCW tilt in Yaw (viewed from the top of the prism). According to the matrix calculation (T1500060) this will induce ~1.5mm shift of the beam. This should be tried before gluing.

  278   Fri May 26 21:53:20 2017 KojiGeneralConfigurationTrans RF PD setup

Recent work

- DC output of the trans RF PD was connected to the BNC patch panel. => Now CH4 of the scope is monitoring this signal

- The RF sweep signal from the network analyzer is connected to the power combiner for the EOM drive via the SMA patch panel.

- The trans RF PD was aligned first to the leakage beam. It turned out that this signal is too weak. Then the PD was aligned to one of the main OMC transmission. For this purpose, the OMC DCPD (T) was removed from the OMC breadboard.

- It seems that there is a significant amount of RF AM from the EOM. I suspect it is associated with the residual S-pol and birefringence of the steering mirrors (45deg HR). But the HWP at the output of the Faraday is fixed on the Faraday body with a screw and cumbersome for fine adjustment. A PBS and an HWP are added right before the EOM. This made the fiber coupler slightly misaligned. I suppose this new setup still has S&P on the fiber too. Thus, readjustment of the fiber rotations at the input is necessary.

Next step

- Input power to the fiber should be determined before the EOM. Otherwise, touching the HWP before the EOM causes too much power change at the optics of the OMC side.

- Precise adjustment of the RFAM is still necessary.

- The OMC curved mirror should be held by the new fixture.

- Check the beam spots

- Measure cavity parameters. (transmission/FSR/HOM/etc)

==> Then the curved mirror and the PZT will be glued on the prism

  277   Tue May 16 19:05:18 2017 KojiOpticsConfigurationOMC SN002 fix - temporary optics

Working on the SN002 OMC fix. Checked the inventory. I think I am using C8 mirror as the new temporary CM1 and PZT24 as the new temporary CM2.

  276   Tue Mar 28 21:04:27 2017 KojiElectronicsCharacterizationPDH amp

Attachment 1: PDH amp RF part (before the preamp was installed)

Attachment 2: RF-AF transmission

Attachment 3: Attachment 3: LO dependence

Attachment 4: RF amp gain (saturation)

Attachment 5: Input/output noise level

Attachment 6: Attachment 6: Preamp/DCPD out buffer AF circuit

Attachment 1: DSC_0269.JPG
DSC_0269.JPG
Attachment 2: RF_to_AF_conversion.pdf
RF_to_AF_conversion.pdf
Attachment 3: LO_dependence.pdf
LO_dependence.pdf
Attachment 4: RFamp_gain.pdf
RFamp_gain.pdf
Attachment 5: PreampNoise.pdf
PreampNoise.pdf
Attachment 6: preamp.png
preamp.png
  275   Thu Feb 16 17:23:12 2017 KojiSupplyGeneralPurchase

 

  274   Thu Jan 19 20:57:53 2017 KojiSupplyGeneralPurchase

Ordered:

Office Depot
v AA battery Qty. 24
v 9V battery Qty. 4
v Floor cable cover (6ft)

Thorlabs
v HV PZT Driver
v Lenses

  273   Thu Dec 8 21:17:09 2016 KojiGeneralGeneralOMC placed on the table / the beam roughly aligned

The OMC cavity was locked. The alignment was precisely adjusted. The mode matching was optimized by the lens positions. The reflection during the lock is ~0.01 compared to the full reflection on non-resonance, meaning the mode matching is ~99%. The error signal was maximized (i.e. demod pahse was adjusted) by sweeping the modulation frequency. Note that the EOM is broad band. The modulation freq chosen today was 34.6MHz.

Some notes:

- The error signal has not been preamplified at all yet. Because of this, the reflection is very much sensitive to the input offset.

- The OMC needs wind shield to prevent from the noise caused by air turbulance.

- The laser PZT was actuated via the Thorlabs HV amp. Otherwise, the thermal path needs to be configured.

- One of the CCD monitor is dead. Needs more replacement.

- All the electronics should be moved to the rack. This required long BNC and SMA cables.

- The optical table needs cleaning.

  272   Wed Dec 7 21:18:35 2016 KojiGeneralGeneralOMC placed on the table / the beam roughly aligned

The OMC mode matching sled was fixed on the nominal part of the table. Then the OMC was located at the nominal position marked by three poles.

The input periscope was adjusted to have the input beam roughtly centered on the OMC QPDs. This made the beam from FM2 aligned to the missing CM1, and the beam just went through the hole of the mounting prism. Very promising!

I wanted to use the new (modified) mirror gluing fixture to hold a curved mirror on the mounting prism. It turned out that the fixture was neither cleaned nor assembled. I will ask Downs Team to help me to get the cleaned and assembled fixtures.

Meanwhile, I just reused the original gluing fixture upside down in order to proceed cavity alignment and locking. (Attachment 1)
In fact, once the mirror is placed on the mounting prism, the cavity started to flash without further alignment. I thank for the very precise (repeatable) alignment of the OMC optics and PD/QPDs.

The next steps are initial cavity locking, more alignment, and mode matching.

Attachment 1: DSC_0082.jpg
DSC_0082.jpg
Attachment 2: DSC_0084.jpg
DSC_0084.jpg
  271   Wed Dec 7 19:18:10 2016 KojiOpticsGeneralLWE NPRO Laser / Input Optics / Fiber Coupling

FIber Input Mount 132deg
Fiber output mount 275deg
-> 525mW P: 517mW S: 8mW extinction ratio: 0.016

  270   Mon Nov 21 21:19:20 2016 KojiOpticsGeneralLWE NPRO Laser / Input Optics / Fiber Coupling

- About 1.5 month ago, an 700mW LWE NPRO has been brought to OMC Lab.

- The SOP can be found here.

- The base was made for the beam elevation of 3" height. Four 1" pedestals were attached to rise the beam elevation to 4".

- The output from the laser is ~740mW

- After the faraday and the BB EOM, the output is ~660mW

- After the usual struggle, the beam was coupled to the SM fiber. The output is 540mW. The coupling efficiency is >80%.

- Will proceed to the OMC cavity alignment.

  269   Fri Sep 9 19:43:32 2016 KojiOpticsGeneralD1102211 OMC Diode Mount Glass Block went to Downs

D1102211 OMC Diode Mount Glass Block (11pcs) have been given to Calum@Downs

  268   Fri Sep 9 14:34:31 2016 KojiGeneralGeneralItem lending

To 40m

First Contact Kit by Calum

Class A Kapton sheets

 

  267   Thu Aug 25 02:17:09 2016 KojiOpticsCharacterizationInspection of the damaged CM1 (prev H1OMC)

Initial inspection results by Calum, et al.
https://dcc.ligo.org/LIGO-E1600268

  266   Tue Aug 23 23:36:54 2016 KojiOpticsCharacterizationInspection of the damaged CM1 (prev H1OMC)

1. Calum and GariLynn checking the CM1 defect from the front side.
2. Same as above
3. Close up of the defect
4. Using dino-lite microscope to get a close up view of the defect from the front surface.
5. Same as 4
6. Finished for the day and setting up a safefy clamp
7. Finally a tefron cover was attached.

Attachment 1: P8238983.jpg
P8238983.jpg
Attachment 2: P8238986.jpg
P8238986.jpg
Attachment 3: P8238987.jpg
P8238987.jpg
Attachment 4: P8238989.jpg
P8238989.jpg
Attachment 5: P8238990.jpg
P8238990.jpg
Attachment 6: P8238994.jpg
P8238994.jpg
Attachment 7: P8238996.jpg
P8238996.jpg
  265   Mon Aug 22 12:58:16 2016 KojiGeneralGeneralUV bond samples -> Garilynn

- FS base + Mounting Prism

- FS or SF2 1/2" piece + FS or SF2 1/2" piece

- FS? plate + FS or SF2 1/2" piece + FS or SF2 1/2" piece + FS? plate

  264   Mon Aug 15 10:09:10 2016 KojiGeneralGeneralPrev H1 OMC shipped to CIT

Previous H1 OMC shipped from LHO to CIT

https://ics-redux.ligo-la.caltech.edu/JIRA/browse/Shipment-8196

  263   Fri Aug 12 14:58:17 2016 KojiGeneralConfigurationH1 OMC DCPD replacement

Preparation of 3rd OMC for the use in H1

New DCPD(T) = B1-01
DCPD(T) = DCPDA: extracted and accomodated in CAGE-A SLOT1

New DCPD(R) = B1-16
DCPD(R) = DCPDB: extracted and accomodated in CAGE-A SLOT2

  262   Fri Jul 22 22:24:05 2016 KojiGeneralGeneralHQEPD inventory

As of Jul 22, 2016
As of Aug 11, 2016

As of Aug 16, 2016


A1-23 in Cage G https://ics-redux.ligo-la.caltech.edu/JIRA/browse/IHGQEX3000-0-00-A1-23
-> Shipped to LLO https://ics-redux.ligo-la.caltech.edu/JIRA/browse/Shipment-8181
-> Now in https://ics-redux.ligo-la.caltech.edu/JIRA/browse/ASSY-D1201439-1
= Replaced C30665 eLIGO PD (SN 01 in Cage G now) ICS: C30665GH-0-00-0001
-> Removed PD@LLO, Waiting for the shipment to CIT

A1-25 in Cage G https://ics-redux.ligo-la.caltech.edu/JIRA/browse/IHGQEX3000-0-00-A1-25
-> Shipped to LLO https://ics-redux.ligo-la.caltech.edu/JIRA/browse/Shipment-8181
-> Now in https://ics-redux.ligo-la.caltech.edu/JIRA/browse/ASSY-D1201439-1
= Replaced C30665 eLIGO PD (SN 02 in Cage G now) ICS: C30665GH-0-00-0002
-> Removed@LLO, Waiting for the shipment to CIT


B1-01 in Cage A https://ics-redux.ligo-la.caltech.edu/JIRA/browse/IHGQEX3000-0-00-B1-01
-> Shipped to LHO https://ics-redux.ligo-la.caltech.edu/JIRA/browse/Shipment-8182
-> Now in https://ics-redux.ligo-la.caltech.edu/JIRA/browse/ASSY-D1201439-3_2
= replaced C30665 eLIGO PD (SN 11 in Cage A now) ICS: C30665GH-0-00-0011
-> Removed PD@LHO
-> Shipped from LHO to CIT https://ics-redux.ligo-la.caltech.edu/JIRA/browse/Shipment-8187

B1-16 in Cage A https://ics-redux.ligo-la.caltech.edu/JIRA/browse/IHGQEX3000-0-00-B1-16
-> Shipped to LHO https://ics-redux.ligo-la.caltech.edu/JIRA/browse/Shipment-8182
-> Now in https://ics-redux.ligo-la.caltech.edu/JIRA/browse/ASSY-D1201439-3_2
= replaced C30665 eLIGO PD (SN 12 in Cage A now) ICS: C30665GH-0-00-0012
-> Removed PD@LHO
-> Shipped from LHO to CIT https://ics-redux.ligo-la.caltech.edu/JIRA/browse/Shipment-8187


C1-05 in Cage F https://ics-redux.ligo-la.caltech.edu/JIRA/browse/IHGQEX3000-0-00-C1-05
-> @CIT contamination test cavity

C1-07 in Cage F https://ics-redux.ligo-la.caltech.edu/JIRA/browse/IHGQEX3000-0-00-C1-07
-> @CIT contamination test cavity


C1-17 in Cage E https://ics-redux.ligo-la.caltech.edu/JIRA/browse/IHGQEX3000-0-00-C1-17
-> Shipped to LHO https://ics-redux.ligo-la.caltech.edu/JIRA/browse/Shipment-8182
-> Left @LHO as a spare

C1-21 in Cage E https://ics-redux.ligo-la.caltech.edu/JIRA/browse/IHGQEX3000-0-00-C1-21
-> Shipped to LHO https://ics-redux.ligo-la.caltech.edu/JIRA/browse/Shipment-8182
-> Left @LHO as a spare


D1-08 in Cage E https://ics-redux.ligo-la.caltech.edu/JIRA/browse/IHGQEX3000-0-00-D1-08
-> Shipped to LHO https://ics-redux.ligo-la.caltech.edu/JIRA/browse/Shipment-8182
-> Moved to Cage A3
-> Shipped from LHO to CIT https://ics-redux.ligo-la.caltech.edu/JIRA/browse/Shipment-8186
-> Arrived at CIT (Aug 16)

D1-10 in Cage E https://ics-redux.ligo-la.caltech.edu/JIRA/browse/IHGQEX3000-0-00-D1-10
-> Shipped to LHO https://ics-redux.ligo-la.caltech.edu/JIRA/browse/Shipment-8182
-> Moved to Cage A4
-> Shipped from LHO to CIT https://ics-redux.ligo-la.caltech.edu/JIRA/browse/Shipment-8186
-> Arrived at CIT (Aug 16)

ELOG V3.1.3-